Shift Summary: Day, 12/14/99

Angela Bellavance (
Tue, 14 Dec 1999 15:43:36 -0600 (CST)

Shift Summary for day shift, 14/Dec/99

On Shift: E. Turner, A. Bellavance

Hours of beam delivered at all intensities: 8.3
Hours of beam delivered at nominal intensity: 8.3
Hours of E799 data (special runs not included): 8.1
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: bad spills, ktevana errors, changing tapes
Runs ended this shift:
Run 15218 - 25M events - BEAM_799

Tapes brought to FCC this shift: UPA217-22

Accesses during shift: none

Overview of shift:
Data taking with many minor alarms (scaler info just slightly out of

Description of Shift:

run spill time comment
15218 7:05 Angela arrives. We are taking good beam.
7:29 OnSplit mount request: UPH040
8:17 OnSplit mount request: UPM044
9:45 High current load alarm on TRD7, one spill only.
10:10 Ktevana errors - KFREND.
314-5 10:11 Beam gone for 2 spills.
316-7 Pipeline problems.
318-20 Beam gone again.
377-9 11:35 Ktevana errors - stream 5, crate 223
382 Crappy spill.
383 Beam gone, 1 spill.
385-6 Trigger hangs. Reset laser repeater power.
12:05 Gas checklist done.
15:00 Run ends with full tapes.
15:10 Onsplit mount request: UPPE09 & UPLA11.
15219 1 15:12 New BEAM799 run started.
15:30 New shift crew is here.

Follow up Notes on Previous Shift

CsI Gain Drift plot did clear up when we started a new run.

Items to Follow up on Next Shift

BobT's project for down time: send Acc.1 upstairs instead of Acc.4

For Elizabeth, Angela

Angela Bellavance work: (630) 840-4144 home: (630) 548-3412 page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke